OpenStudy (anonymous):

log 256 b greater than 1/4

OpenStudy (anonymous):

answers are b greater than 1/4 b greater than 4 b greater than 1/4 and b greater than 2^3

OpenStudy (anonymous):

b greater than 1/256

OpenStudy (jdoe0001):

$$\bf log_{256}b = \cfrac{1}{4}\\ \textit{using log cancellation rule of } log_aa^x = x\\ 256^{log_{256}b} = 256^{\frac{1}{4}} \implies b = 256^{\frac{1}{4}} \implies b = \sqrt[4]{256} $$

OpenStudy (anonymous):

but with what answer will it correspond with

OpenStudy (jdoe0001):

well, get the value of b

OpenStudy (anonymous):

the value is 4

OpenStudy (jdoe0001):

then that's "b" :)